OpenStudy (anonymous):

CALC QUESTION: (Will medal) An object moves along the x-axis with velocity of v(t) = t(sqrt(2t+3)) -3t Find the acceleration of a(t) of the object. I understand I need to find the derivative. But having issues

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

\[t \sqrt{2t+3} - 3t\]

OpenStudy (misty1212):

HI!!you have a product, so you need the product rule for the first part

OpenStudy (anonymous):

Hi!

OpenStudy (misty1212):

the derivative of \(-3t\) is easy right? it is just \(-3\)

OpenStudy (misty1212):

\[\left(fg\right).=f'g+g'f\] with \[f(t)=t, f'(t)=1,g(t)=\sqrt{2t+3}, g'(t)=\frac{1}{\sqrt{2t+3}}\]

OpenStudy (anonymous):

Yes haha! Simple as that?

OpenStudy (anonymous):

Then will we have to use chain rule?

OpenStudy (misty1212):

to find the derivative o f\(\sqrt{2x+3}\) you use the chain rule (in your head)

OpenStudy (misty1212):

the derivative of the square root of sommat is one over two square root of sommat, times the derivative of sommat in this case the derivative of \(2x+3\) is \(2\) so the twos cancel

OpenStudy (janu16):

hey misty do you think you can help me after youre done with this?

OpenStudy (anonymous):

Ah ok! Thank you for your time!

OpenStudy (anonymous):

Just confirming.... (sorry) @misty1212, I got the answer to be t - 3. Is that right?